Pearce & Company has 10 million shares of $2 par value common stock outstanding. The company believes that its current market price of $100 per share is too high and decides to execute a 4-for-1 forward stock split to lower the price. How many shares will be outstanding following the stock split, and what will be the new par value per share?
Answer #1

Number f shares = 10,000,000 *4/1 = 40,000,000 shares

par value = 2 *1/4 = $ .50 per share